If x = e^ty^2 , t= V cos v , y= v sin v. Find dx/dv at v= 3.142/2

1
Expert's answer
2021-11-23T09:47:05-0500
"x=e^ty^2"

"x(v)=e^{v\cos(v)}v^2\sin^2(v)"

"dx/dv=e^{v\cos(v)}(\cos(v)-v\sin(v))"

"+e^{v\cos(v)}(2v\sin(v)+2v^2\sin(v)\cos(v))"

"v=\pi/2"


"dx/dv|_{v=\pi/2}=e^{0}(0-\pi/2)"

"+e^{0}(2(\pi/2)(1)+2(\pi/2)^2(1)(0)) =-\pi/2+\pi=\pi/2"